# Spectroastrometry and Reverberation Mapping of Active Galactic Nuclei. II. Measuring Geometric Distances and Black Hole Masses of Four Nearby Quasars

### v1

@@ Abstract
The geometric distances of active galactic nuclei (AGNs) are challenging to measure because of their exceptionally compact structure yet vast cosmic distances. A combination of spectroastrometry and reverberation mapping (SARM) of broad-line regions (BLRs) constitutes a novel means to probe the geometric distance of AGNs, which has recently become practically feasible owing to successful interferometric observations with VLTI/GRAVITY. Here, we perform SARM analysis of four nearby quasars: Mrk 509, PDS 456, 3C 273, and NGC 3783. Results for the former two are reported for the first time and the latter two are revisited using our improved BLR dynamical modeling that includes the radial-dependent responsivity of BLRs. This allows us to self-consistently account for the emissivity weighting of the BLR in spectroastrometry and responsivity weighting in reverberation mapping. We obtain angular-diameter distances of the four quasars, from which we derive a Hubble constant of . Although this consititutes a large uncertainty for a measurement of , it is anticipated that the precision will improve to a competitive level once a greater number of AGNs are accessible following the upgrade of GRAVITY in the near future. From SARM analysis, the black hole masses of the four quasars are also measured with the statistical uncertainty ranging from 0.06 to 0.23 dex, consistent with the correlations between black hole masses and properties of the host bulges.
@@ 1 Introduction
Over the past decades, major efforts have been made to spectroscopically monitor AGNs continually over long periods, enabling measurement of the BLR sizes through RM analysis. The typical sizes of BLRs were found to range from light-days to light-months, mainly depending on the AGN luminosity (e.g., Kaspi et al. 2000; Bentz et al. 2013; Du & Wang 2019). Such sizes are too compact to be spatially resolved by modern facilities even for the nearest AGNs. For instance, the Seyfert 1 galaxy NGC 4151 at a distance of 15.8 Mpc (Yuan et al. 2020) has a typical BLR size of light-days (Chen et al. 2023), corresponding to an angular size of milli-arcseconds. For a long time RM stood out as the most effective method to determine the spatial extent of the BLR (e.g., Kaspi et al. 2000; Peterson et al. 2004; Bentz et al. 2009; Du et al. 2014). It has also been an essential tool for the diagnostic of the gas environment in the vicinity of SMBHs (Cackett et al. 2021).
@@
Remarkably, the capabilities of SA and RM techniques are complementary and their combination (hereafter referred to as SARM) conveys a more comprehensive view of the BLR geometry and kinematics. This is reflected in the different dimensions probed by SA and RM. To be specific, SA is sensitive to the BLR structure projected onto the observer’s sky, i.e., perpendicular to the line of sight. In contrast, RM is sensitive to the structure perpendicular to the iso-delay surfaces since it measures time delays between variations of the emission line and continuum. The iso-delay surfaces have a paraboloid shape with the focus located at the central ionizing source, typically regarded as a point source coincident with the SMBH. The combination of the results from both techniques thus enable a much better overall understanding of the BLR properties as well as precise SMBH mass measurements.
@@
Another important application of the SARM method is that it directly probes geometric distances to AGNs. The time delays measured from RM represent light travel times from the central ionizing source, typically an accretion disk, to the BLR. Multiplied by the speed of light, these thereby convey the linear size () scale of the BLR. The interferometric phases measured from SA reflect the angular astrometry of BLR photons and convey the angular size scale () of the BLR. A combination of SA and RM naturally constitutes a geometric probe for the angular-diameter distance () from the basic geometric principles (Elvis & Karovska 2002; Rakshit et al. 2015; Wang et al. 2020; Li & Wang 2023). As it is usually difficult to measure both and for the same target, extragalactic objects for which such a geometric method is applicable are exceptionally rare. The SARM approach thus offers a new, promising tool for cosmology, especially given ubiquitousness and high brightness of AGNs across the universe.
@@
Hitherto a sample of seven nearby bright AGNs has been observed by GRAVITY (GRAVITY Collaboration et al. 2018, 2020, 2021a, 2024). Wang et al. (2020, hereafter W20) conducted the first SARM analysis on 3C 273 and GRAVITY Collaboration et al. (2021b, hereafter G21b) conducted the second application of the SARM technique on NGC 3783. These initial studies derived an angular-diameter distance Mpc with a precision of 15% for 3C 273, and Mpc with a precision of 30% for NGC 3783, respectively. They adopted a globally linear, uniform response of the BLR gas. However, there is evidence that the BLR gas might have a radial-dependent, non-linear response. The most relevant observational phenomenon comes from RM campaigns, which show that the H line widths of the mean spectra are typically broader than those of the root-mean-square (rms) spectra (see Figure 6 in Li & Wang 2025). This clearly implies that the BLR gas is more responsive at large radial distances, where the BLR gas has a relatively lower velocity (compared to the inner regions), thereby resulting in a narrower line width in the rms spectrum. Such a non-uniform response of BLRs has also been shown from the photoionization calculations (e.g., Korista & Goad 2004; Goad & Korista 2014; Zhang et al. 2021). If using the terminology “responsivity”, which quantifies the relative change of the line emissivity with respect to that of the incident continuum flux (e.g., Goad et al. 1993), the responsivity of the BLR gas is not constant, but changes with radial distance. Based on photoionization theory, the main physical reason is as follows. In the inner region where BLR gas receives a higher ionizing photon flux, the excited-state populations of hydrogen and helium atoms are enhanced, leading to a lower efficiency in converting the ionizing photons into recombination lines (such as the Balmer lines) that arise out of these excited states (Korista & Goad 2004). As a consequence, the BLR gas exhibits a low responsivity. In an extreme case where the BLR gas is fully ionized, the emission lines will no longer respond to any continuum variations, resulting in a zero responsivity. As the distance () to the ionizing source increases, the incident ionizing photon flux decreases () and the above effect gradually weakens. In other words, for an extended BLR, the changes in the responsivity of emission lines with respect to the incident ionizing photon flux naturally cause a dependence of the responsivity on radial distance.
@@
This radially dependent responsivity of BLR gas has an important implication for SARM analysis. As SA probes emissivity weighting whereas RM probes responsivity weighting of the BLR gas, a non-uniform responsivity means that the emissivity- and responsivity-weighted BLR sizes are not the same. In this work, we perform SARM analysis on two AGNs investigated by GRAVITY Collaboration et al. (2024), Mrk 509 and PDS 456, and revisit the previous analysis on 3C 273 and NGC 3783 using the latest improved BLR dynamical modeling that includes radially dependent responsivity of the BLR gas (Li & Wang 2025). This allows us to self-consistently deal with emissivity weighting in SA and responsivity weighting in RM and therefore to alleviate inaccuracies arising from the adoption of globally uniform responses in previous analysis.
@@
The paper is organized as follows. Section 2 compiles the observation data of SA from VLTI/GRAVITY and RM from previous optical spectral monitoring campaigns. Section 3 introduces the methodology of SARM joint analysis, including continuum light curve modeling, BLR dynamical modeling, and Bayesian inference. Section 4 summarizes estimates of the geometric distances, Hubble constant, and the black hole masses. Section 5 presents discussions on possible systematic uncertainties in the determined distances, followed by the conclusions in Section 6. Appendix A defines three types of BLR sizes, Appendix B lists the full set of model parameters and their inferred values, Appendix C shows the best fits to the differential phases of the three baselines with the strongest BLR signals, Appendix D plots the mean and rms spectra of the four AGNs, and Appendix E tests the influence of the upper limit of responsivity on the inferred model parameters.
@@
GRAVITY Period | Jul. 2021 (2 nights) | Aug. 2018-Jul. 2021 (9 nights) | Jul. 2017-May 2018 (8 nights) | Jan. 2018-Mar. 2020 (6 nights) |
RM Period | May 2019-Nov. 2022 | May 2020-Oct. 2023 | Nov. 2008-Jul. 2018 | Feb. 2020-Jun. 2020 |
@@ 2.1 GRAVITY Interferometric Data
We made additional reduction to the differential phase and line profile data of Mrk 509 and PDS 456 after applying the pipeline following the procedures described in GRAVITY Collaboration et al. (2020). We first fit and remove the static and variable instrumental effect of the differential phase for individual exposures. We also use a first-order polynomial function to remove the residual phase at the continuum emission wavelength (see Section 2.3 of GRAVITY Collaboration et al. 2020). We determine the phase error from the standard deviation of the phase at the continuum for each exposure and average all the exposures weighted according to their phase errors. To calibrate the broad Br or Pa line, we used the calibrator stars observed before or after the AGN observations to remove the instrumental and telluric features. Thereafter, we apply a third-order polynomial function to fit the calibrated continuum and normalize the line profile to the continuum. For Mrk 509 and PDS 456, we prefer using early-type stars to calibrate the spectra because they have less absorption features, with only the Br absorption line significant in the relevant spectral range, which imprints a narrow Br emission line on the final calibrated AGN spectra. To subtract this contaminating component, we use two Gaussian functions to fit the AGN Br or Pa line and a Lorentzian function to fit the narrow Br line. The AGN and stellar Br lines are not strongly blended (if at all) because the AGN Br line is significantly redshifted. We notice that, compared to the double-Gaussian function used to fit the stellar Br line as by GRAVITY Collaboration et al. (2024), the Lorentzian function can improve the fit to the Pa line profile of PDS 456. This is because the double-Gaussian function for the stellar Br line cannot be properly separated from those for the AGN Pa line, thereby causing an inappropriate subtraction around the blue wing of the AGN line. We directly adopt the data of 3C 273 and NGC 3783 from GRAVITY Collaboration et al. (2018) and GRAVITY Collaboration et al. (2021a), respectively. The NGC 3783 data were reduced in the same way as described above. The 3C 273 data were however reduced differently. Instead of decomposing the instrumental differential phase signatures into static and variable components, GRAVITY Collaboration et al. (2018) incorporated a Gaussian kernel with an FWHM of 24 pixels to convolve the differential phase spectra and generate the broad instrumental differential phase. We confirmed that it makes little difference if we conduct the joint analysis with the data reduced with the new approach.
@@ 2.2 Reverberation Mapping Data
Li et al. (2024) reported multi-year RM data of five AGNs (IC 4329A, Mrk 335, Mrk 509, Mrk 1239, and PDS 456), selected from the sample of potential AGNs suited for GRAVITY/GRAVITY+ observations compiled by W20. For Mrk 509 and PDS 456, the optical spectroscopic and photometric monitoring were conducted from May 2019 to November 2022, and from May 2020 to October 2023, respectively, with a typical sampling cadence of every 4-5 days. The photometric data were further combined with the public archival data from the Zwicky Transient Facility (ZTF; Graham et al. 2019) and the All-Sky Automated Survey for Supernovae (ASAS-SN; Shappee et al. 2014; Kochanek et al. 2017) to yield high-cadence continuum light curves. These different sources of data were intercalibrated with the Bayesian software PyCALI111https://github.com/LiyrAstroph/%****␣ms_v5.tex␣Line␣225␣****PyCALI (Li et al. 2014), which uses a multiplicative factor and an additive factor to account for the inhomogeneous apertures and differences in flux calibration. For Mrk 509, from the four years of monitoring, the H time delay (around 50 days) can be detected in three years (2019, 2021, and 2022). However, when using the full data span, there appears a long-term trend in the H variations, leading to an apparently long time delay component (around 300 days). The explanation for such a long-term trend is still under debate. We therefore only use one-year data for SARM analysis. We choose the RM data for year 2022 since it shows the most significant variability. We also present analysis with the RM data in 2019 and 2021 and find that the obtained distances are consistent with each other within uncertainties. For PDS 456, the H time delay is about 250 days and we use the four-year RM data for the SARM analysis.
@@
The latest published RM study of NGC 3783 was conducted by Bentz et al. (2021) throughout the first half of 2020 using the Las Cumbres Observatory global telescope network. A total of 50 spectra were obtained with a typical sampling interval of 2 days. The 5100 Å data was merged with the -band photometric data (with 209 epochs), resulting in a continuum light curve spanning a time span of 137 days from JD 2,458,891 to 2,459,028 (see Figure 4 in Bentz et al. 2021). NGC 3783 has also been covered by the ASAS-SN survey since February 2012. We use the PyCALI software to scale and shift the ASAS-SN light curve to match the merged light curve of Bentz et al. (2021). This extends the temporal baseline of the continuum light curve from 137 days to 268 days in the 2020 season and also supplies continuum variation information for other seasons. The former is useful to constrain the BLR model and the latter helps to better characterize the continuum variability.
In modeling RM data, we need to subtract the components in the continuum light curves irrelevant to emission-line reverberation. One such component is the host galaxy’s starlight. Among the four AGNs, Mrk 509, PDS 456, and 3C 273 have a high optical luminosity at 5100 Å (Zhang et al. 2019; Li et al. 2024) and the host galaxy contribution can be treated as negligible. This is supported by the spectral decomposition, which did not identify a significant host galaxy component (Zhang et al. 2019; Li et al. 2024). For NGC 3783, we subtract the host galaxy component () according to the HST image decomposition employed by Bentz et al. (2021). 3C 273 is additionally subject to the contamination from the jet emission. We subtract an averaged jet flux from the continuum light curve according to the decomposition of the jet and accretion disk emissions proposed by Li et al. (2020), which derived an averaged jet flux to the total continuum flux fraction of .

### v2

@@ Abstract
The geometric distances of active galactic nuclei (AGNs) are challenging to measure because of their exceptionally compact structure yet vast cosmic distances. A combination of spectroastrometry and reverberation mapping (SARM) of broad-line regions (BLRs) constitutes a novel means to probe the geometric distance of AGNs, which has recently become practically feasible owing to successful interferometric observations with VLTI/GRAVITY. Here, we perform SARM analysis of four nearby quasars: Mrk 509, PDS 456, 3C 273, and NGC 3783. Results for the former two are reported for the first time and the latter two are revisited using our improved BLR dynamical modeling that includes the radial-dependent responsivity of BLRs. This allows us to self-consistently account for the emissivity weighting of the BLR in spectroastrometry and responsivity weighting in reverberation mapping. We obtain angular-diameter distances of the four quasars, from which we derive a Hubble constant of . Although this constitutes a large uncertainty for a measurement of , it is anticipated that the precision will improve to a competitive level once a greater number of AGNs are accessible following the upgrade of GRAVITY in the near future. From SARM analysis, the black hole masses of the four quasars are also measured with the statistical uncertainty ranging from 0.06 to 0.23 dex, consistent with the correlations between black hole masses and properties of the host bulges.
@@ 1 Introduction
Over the past decades, major efforts have been made to spectroscopically monitor AGNs continually over long periods, enabling measurement of the BLR sizes through RM analysis. The typical sizes of BLRs were found to range from light-days to light-months, mainly depending on the AGN luminosity (e.g., Kaspi et al. 2000; Bentz et al. 2013; Du & Wang 2019). Such sizes are too compact to be spatially resolved by modern facilities, even for the nearest AGNs. For instance, the Seyfert 1 galaxy NGC 4151 at a distance of 15.8 Mpc (Yuan et al. 2020) has a typical BLR size of light-days (Chen et al. 2023), corresponding to an angular size of milli-arcseconds. For a long time, RM stood out as the most effective method to determine the spatial extent of the BLR (e.g., Kaspi et al. 2000; Peterson et al. 2004; Bentz et al. 2009; Du et al. 2014). It has also been an essential tool for the diagnostic of the gas environment in the vicinity of SMBHs (Cackett et al. 2021).
@@
Remarkably, the capabilities of SA and RM techniques are complementary and their combination (hereafter referred to as spectroastrometry and reverberation mapping (SARM)) conveys a more comprehensive view of the BLR geometry and kinematics. This is reflected in the different dimensions probed by SA and RM. To be specific, SA is sensitive to the BLR structure projected onto the observer’s sky, i.e., perpendicular to the line of sight. In contrast, RM is sensitive to the structure perpendicular to the iso-delay surfaces since it measures time delays between variations of the emission line and continuum. The iso-delay surfaces have a paraboloid shape with the focus located at the central ionizing source, typically regarded as a point source coincident with the SMBH. The combination of the results from both techniques thus enables a much better overall understanding of the BLR properties as well as precise SMBH mass measurements.
@@
Another important application of the SARM method is that it directly probes geometric distances to AGNs. The time delays measured from RM represent light travel times from the central ionizing source, typically an accretion disk, to the BLR. Multiplied by the speed of light, these thereby convey the linear size () scale of the BLR. The interferometric phases measured from SA reflect the angular astrometry of BLR photons and convey the angular size scale () of the BLR. A combination of SA and RM naturally constitutes a geometric probe for the angular-diameter distance () from the basic geometric principles (Elvis & Karovska 2002; Rakshit et al. 2015; Wang et al. 2020; Li & Wang 2023). As it is usually difficult to measure both and for the same target, extragalactic objects for which such a geometric method is applicable are exceptionally rare. The SARM approach thus offers a new, promising tool for cosmology, especially given the ubiquitousness and high brightness of AGNs across the Universe.
@@
Hitherto, a sample of seven nearby bright AGNs has been observed by GRAVITY (GRAVITY Collaboration et al. 2018, 2020, 2021a, 2024). Wang et al. (2020, hereafter W20) conducted the first SARM analysis on 3C 273 and GRAVITY Collaboration et al. (2021b, hereafter G21b) conducted the second application of the SARM technique on NGC 3783. These initial studies derived an angular-diameter distance Mpc with a precision of 15% for 3C 273, and Mpc with a precision of 30% for NGC 3783, respectively. They adopted a globally linear, uniform response of the BLR gas. However, there is evidence that the BLR gas might have a radial-dependent, non-linear response. The most relevant observational phenomenon comes from RM campaigns, which show that the H line widths of the mean spectra are typically broader than those of the root-mean-square (rms) spectra (see Figure 6 in Li & Wang 2025). This clearly implies that the BLR gas is more responsive at large radial distances, where the BLR gas has a relatively lower velocity (compared to the inner regions), thereby resulting in a narrower line width in the rms spectrum. Such a non-uniform response of BLRs has also been shown from the photoionization calculations (e.g., Korista & Goad 2004; Goad & Korista 2014; Zhang et al. 2021). If using the terminology “responsivity”, which quantifies the relative change of the line emissivity with respect to that of the incident continuum flux (e.g., Goad et al. 1993), the responsivity of the BLR gas is not constant, but changes with radial distance. Based on photoionization theory, the main physical reason is as follows. In the inner region where BLR gas receives a higher ionizing photon flux, the excited-state populations of hydrogen and helium atoms are enhanced, leading to a lower efficiency in converting the ionizing photons into recombination lines (such as the Balmer lines) that arise out of these excited states (Korista & Goad 2004). As a consequence, the BLR gas exhibits a low responsivity. In an extreme case where the BLR gas is fully ionized, the emission lines will no longer respond to any continuum variations, resulting in a zero responsivity. As the distance () to the ionizing source increases, the incident ionizing photon flux decreases () and the above effect gradually weakens. In other words, for an extended BLR, the changes in the responsivity of emission lines with respect to the incident ionizing photon flux naturally cause a dependence of the responsivity on radial distance.
@@
This radially dependent responsivity of BLR gas has an important implication for SARM analysis. As SA probes emissivity weighting, whereas RM probes responsivity weighting of the BLR gas, a non-uniform responsivity means that the emissivity- and responsivity-weighted BLR sizes are not the same. In this work, we perform SARM analysis on two AGNs investigated by GRAVITY Collaboration et al. (2024), Mrk 509 and PDS 456, and revisit the previous analysis on 3C 273 and NGC 3783 using the latest improved BLR dynamical modeling that includes radially dependent responsivity of the BLR gas (Li & Wang 2025). This allows us to self-consistently deal with emissivity weighting in SA and responsivity weighting in RM and therefore to alleviate inaccuracies arising from the adoption of globally uniform responses in previous analysis.
@@
The paper is organized as follows. Section 2 compiles the observation data of SA from VLTI/GRAVITY and RM from previous optical spectral monitoring campaigns. Section 3 introduces the methodology of SARM joint analysis, including continuum light curve modeling, BLR dynamical modeling, and Bayesian inference. Section 4 summarizes estimates of the geometric distances, Hubble constant, and the black hole masses. Section 5 presents discussions on possible systematic uncertainties in the determined distances, followed by the conclusions in Section 6. Appendix A defines three types of BLR sizes, Appendix B lists the full set of model parameters and their inferred values, Appendix C shows the best fits to the differential phases of the three baselines with the strongest BLR signals, Appendix D plots the mean and rms spectra of the four AGNs, Appendix E tests the influence of the upper limit of responsivity on the inferred model parameters, and Append F compares the obtained time lags from BLR dynamical modeling with those derived from the traditional cross-correlation analysis of the RM data.
@@
GRAVITY | Jul. 2021 (2 nights) | Aug. 2018-Jul. 2021 (9 nights) | Jul. 2017-May 2018 (8 nights) | Jan. 2018-Mar. 2020 (6 nights) |
RM | May 2019-Nov. 2022 | May 2020-Oct. 2023 | Nov. 2008-Jul. 2018 | Feb. 2020-Jun. 2020 |
@@ 2.1 GRAVITY Interferometric Data
We made additional reductions to the differential phase and line profile data of Mrk 509 and PDS 456 after applying the pipeline following the procedures described in GRAVITY Collaboration et al. (2020). We first fit and remove the static and variable instrumental effect of the differential phase for individual exposures. We also use a first-order polynomial function to remove the residual phase at the continuum emission wavelength (see Section 2.3 of GRAVITY Collaboration et al. 2020). We determine the phase error from the standard deviation of the phase at the continuum for each exposure and average all the exposures weighted according to their phase errors. To calibrate the broad Br or Pa line, we used the calibrator stars observed before or after the AGN observations to remove the instrumental and telluric features. Thereafter, we apply a third-order polynomial function to fit the calibrated continuum and normalize the line profile to the continuum. For Mrk 509 and PDS 456, we prefer using early-type stars to calibrate the spectra because they have less absorption features, with only the Br absorption line significant in the relevant spectral range, which imprints a narrow Br emission line on the final calibrated AGN spectra. To subtract this contaminating component, we use two Gaussian functions to fit the AGN Br or Pa line and a Lorentzian function to fit the narrow Br line. The AGN and stellar Br lines are not strongly blended (if at all) because the AGN Br line is significantly redshifted. We notice that, compared to the double-Gaussian function used to fit the stellar Br line as by GRAVITY Collaboration et al. (2024), the Lorentzian function can improve the fit to the Pa line profile of PDS 456. This is because the double-Gaussian function for the stellar Br line cannot be properly separated from those for the AGN Pa line, thereby causing an inappropriate subtraction around the blue wing of the AGN line. We directly adopt the data of 3C 273 and NGC 3783 from GRAVITY Collaboration et al. (2018) and GRAVITY Collaboration et al. (2021a), respectively. The NGC 3783 data were reduced in the same way as described above. The 3C 273 data were, however, reduced differently. Instead of decomposing the instrumental differential phase signatures into static and variable components, GRAVITY Collaboration et al. (2018) incorporated a Gaussian kernel with an FWHM of 24 pixels to convolve the differential phase spectra and generate the broad instrumental differential phase. We confirmed that it makes little difference if we conduct the joint analysis with the data reduced with the new approach.
@@ 2.2 Reverberation Mapping Data
Li et al. (2024) reported multi-year RM data of five AGNs (IC 4329A, Mrk 335, Mrk 509, Mrk 1239, and PDS 456), selected from the sample of potential AGNs suited for GRAVITY/GRAVITY+ observations compiled by W20. For Mrk 509 and PDS 456, the optical spectroscopic and photometric monitoring were conducted from May 2019 to November 2022, and from May 2020 to October 2023, respectively, with a typical sampling cadence of every 4-5 days. The photometric data were further combined with the public archival data from the Zwicky Transient Facility (ZTF; Graham et al. 2019) and the All-Sky Automated Survey for Supernovae (ASAS-SN; Shappee et al. 2014; Kochanek et al. 2017) to yield high-cadence continuum light curves. These different sources of data were intercalibrated with the Bayesian software PyCALI111The latest code is publicly available at https://github.com/LiyrAstroph/%****␣ms_v5.tex␣Line␣225␣****PyCALI, while the version used in this work is available in Li (2024b). (Li et al. 2014), which uses a multiplicative factor and an additive factor to account for the inhomogeneous apertures and differences in flux calibration. For Mrk 509, from the four years of monitoring, the H time delay (around 50 days) can be detected in three years (2019, 2021, and 2022). However, when using the full data span, there appears to be a long-term trend in the H variations, leading to an apparently long time delay component (around 300 days). The explanation for such a long-term trend is still under debate. We therefore only use one-year data for SARM analysis. We chose the RM data for the year 2022 since it shows the most significant variability. We also present analysis with the RM data in 2019 and 2021 and find that the obtained distances are consistent with each other within uncertainties. For PDS 456, the H time delay is about 250 days and we use the four-year RM data for the SARM analysis.
@@
The latest published RM study of NGC 3783 was conducted by Bentz et al. (2021) throughout the first half of 2020 using the Las Cumbres Observatory global telescope network. A total of 50 spectra were obtained with a typical sampling interval of 2 days. The 5100 Å data was merged with the -band photometric data (with 209 epochs), resulting in a continuum light curve spanning a time span of 137 days from JD 2,458,891 to 2,459,028 (see Figure 4 in Bentz et al. 2021). NGC 3783 has also been covered by the ASAS-SN survey since February 2012. We use the PyCALI software to scale and shift the ASAS-SN light curve to match the merged light curve of Bentz et al. (2021). This extends the temporal baseline of the continuum light curve from 137 days to 268 days in the 2020 season and also supplies continuum variation information for other seasons. The former is useful to constrain the BLR model, and the latter helps to better characterize the continuum variability.
In modeling RM data, we need to subtract the components in the continuum light curves irrelevant to emission-line reverberation. One such component is the host galaxy’s starlight. Among the four AGNs, Mrk 509, PDS 456, and 3C 273 have a high optical luminosity at 5100 Å (Zhang et al. 2019; Li et al. 2024) and the host galaxy contribution can be treated as negligible. This is supported by the spectral decomposition, which did not identify a significant host galaxy component (Zhang et al. 2019; Li et al. 2024). For NGC 3783, we subtract the host galaxy component () according to the Hubble Space Telescope image decomposition employed by Bentz et al. (2021). 3C 273 is additionally subject to the contamination from the jet emission. We subtract an averaged jet flux from the continuum light curve according to the decomposition of the jet and accretion disk emissions proposed by Li et al. (2020), who derived an averaged jet flux to the total continuum flux fraction of .